OpenStudy (anonymous):

y=3x-24 5x-y=8 Walk me through?

jigglypuff314 (jigglypuff314):

if you would like to use the substitution method you first isolate the y in the second equation then y=y so would then substitute what the ys equals (the parts with the x in them) then you solve for x and once you've got that, you can plug the x you got back into one of the original equations

OpenStudy (anonymous):

How do that ?

jigglypuff314 (jigglypuff314):

to isolate the y in the second equation of 5x - y = 8 you minus y on both sides to get 5x = 8 + y then you minus 8 on both sides to get y = 5x - 8 then y = y so 3x - 24 = 5x - 8

OpenStudy (anonymous):

thank you so much